OpenStudy (anonymous):

trying to solve these word problems suppose that the smiths are expecting a total of 24 guests for easter dinner. each guest will get a box of peeps, 3 hot cross buns, and 1/3 of a pound of ham. the total cost of peeps, buns and ham for all the guests is $92.30. A dozen hot cross buns costs 56 cents more than a box of peeps and a pound of ham costs twice as much as the combined cost of a box of peeps and a dozen hot cross buns. find the cost of a box of peeps.

OpenStudy (dumbcow):

let p,b,h represent cost of peeps, buns,and ham 24p + 72b +8h = 92.30 12b = p + 0.56 h = 2(p+12b) solve system with substitution --> h = 2(p+p+.56) = 2(2p+.56) = 4p+1.12 --> b = p/12 + .56/12 24p + (6p+3.36) +(32p+8.96) = 92.30 62p + 12.32 = 92.30 62p = 79.98 p = 1.29

OpenStudy (anonymous):

why is h=2(p + p ?

OpenStudy (dumbcow):

h = 2(p+12b) and 12b = p+.56 i replaced 12b with p+.56

OpenStudy (anonymous):

thank you